OpenStudy (anonymous):

2. Write the equation of the line which passes through (2, 1) and is perpendicular to x = –2.

OpenStudy (anonymous):

How would you write it in standard form?

OpenStudy (compassionate):

\(\Huge{\color{purple}{\textbf{W}} \color{orange}{\cal{E}} \color{green}{\mathbb{L}} \color{blue}{\mathsf{C}} \color{maroon}{\rm{O}} \color{red}{\tt{M}} \color{gold}{\tt{E}} \space \color{orchid}{\mathbf{T}} \color{Navy}{\mathsf{O}} \space \color{OrangeRed}{\boldsymbol{O}} \color{Olive}{\mathbf{P}} \color{Lime}{\textbf{E}} \color{DarkOrchid}{\mathsf{N}} \color{Tan}{\mathtt{S}} \color{magenta}{\mathbb{T}} \color{goldenrod}{\mathsf{U}} \color{ForestGreen}{\textbf{D}} \color{Salmon}{\mathsf{Y}} \ddot \smile }\) The line is perpendicular, meaning it'll have the opposite slope. We are given the two points: (2, 1) Along with the equation: x = -2 Slope: 2 Points: (2, 1) Now yo want to plug it into the formula: \[y - y1 = m(x - x1)\] Where: x1 = 2 y1 = 1 And m = 2

OpenStudy (anonymous):

@Compassionate won't the slope of the line be 1/2??

OpenStudy (compassionate):

2 is the same as 1/2

OpenStudy (anonymous):

Yeah, I got 1/2 as the slope but I got it wrong so now I'm fixing it

terenzreignz (terenzreignz):

1/2 ? 2 ? Guys, the slope of the line x = -2 (this is a vertical line) is undefined, meaning, for a line to be perpendicular to this, the line should have a slope 0...

OpenStudy (anonymous):

oh okay..... sorry for the mistake then..

terenzreignz (terenzreignz):

So, we're looking for a horizontal line which passes through the point (2,1)

OpenStudy (compassionate):

Pardon me, but the slope is zero for any vertical line.

terenzreignz (terenzreignz):

Uhh, no, vertical lines don't have a slope zero, they have an undefined slope, an infinite slope, if you will...

OpenStudy (anonymous):

Compassionate, vertical lines have an undefined slope because: \[m = \frac{dy}{dx}\] So if the change in y is large and the change in x is small then we have an indeterminate quantity. Horizontal lines have a slope of 0 because delta y is zero for any finite delta x.